Suppose that the effective 6-month interest rate is 4.0 percent in the United States and the effective 6-month interest rate in Germany is 8 percent, and that the spot exchange rate is 1.60 USD/EUR and the forward exchange rate, with six-month maturity, is 1.58 USD/EUR.

A. Clearly show whether IRP condition holds or not and explain whether there is an arbitrage opportunity for the home or the foreign investor or neither.

B. Assume that an arbitrageur can borrow up to $1,000,000 or €625,000. What is the profit or loss (in USD) from engaging in covered interest arbitrage? Clearly show the steps involved in your arbitrage strategy.
Answer #1

Fair forward rate as per Interest Rate parity = Spot rate*(1+Interest rate USD)/(1+Interest rate EUR)

= 1.60*(1+4%*6/12)/(1+8%*6/12)

= USD 1.5692/EUR

Since Actual rate is different, arbitrage is possible

B. Borrow $1,000,000

Convert into EUR at spot rate and get 1,000,000/1.60 = EUR 625,000

Invest for 6 months and get 625000(1+8%*6/12) = Eur 650,000

Convert into USD at forward rate = 650,000*1.58 = USD 1,027,000

Repay Loan = 1,000,000*(1+4%*6/12) = 1,020,000

Arbitrage profit = $7,000